A baseball player's batting average is found by dividing the number of hits the player has by the number of at-bats the player has. Suppose a baseball player has 45 hits and 130 at-bats. Write and solve an equation to model the number of consecutive hits the player needs in order to raise his batting average to 0.400. Explain how you found your answer.
Answer by Boreal(15235)

x= number of hits needed, Each hit equals an at bat.
(45+x)/(130+x) =0.4
multiply both sides by (130+x)
45+x=52+ 0.4 x
0.6 x=7; divide by 0.6.
x=12 (11.667)
check 57/142 (12 hits) 0.401
56/141 (11 hits) 0.397
So 12 hits are necessary, since there are no fractions of hits.